# A Majorization-Minimization Gauss-Newton Method for 1-Bit Matrix Completion

### v2

@@ Abstract
In 1-bit matrix completion, the aim is to estimate an underlying low-rank matrix from a partial set of binary observations. We propose a novel method for 1-bit matrix completion called MMGN. Our method is based on the majorization-minimization (MM) principle, which converts the original optimization problem into a sequence of standard low-rank matrix completion problems. We solve each of these sub-problems by a factorization approach that explicitly enforces the assumed low-rank structure and then apply a Gauss-Newton method. Using simulations and a real data example, we illustrate that in comparison to existing 1-bit matrix completion methods, MMGN outputs comparable if not more accurate estimates. In addition, it is often significantly faster, and less sensitive to the spikiness of the underlying matrix. In comparison with three standard generic optimization approaches that directly minimize the original objective, MMGN also exhibits a clear computational advantage, especially when the fraction of observed entries is small.
@@ 1 Introduction
In this work, we consider the following 1-bit matrix completion problem: Let ${\bm{\mathbf{{\Theta}}}}^{*}\in\mathbb{R}^{m\times n}$ be an unknown matrix of exact rank $r^{*}$. The observed data are a subset $\Omega\subset[m]\times[n]$ and a binary matrix ${\bm{\mathbf{{Y}}}}$ whose entries are given by $y_{ij}=0$ if $(i,j)\notin\Omega$, whereas the observed entries are independent draws as follows,
@@
where ${\theta}_{ij}^{*}$ is the $ij$-th entry of ${\bm{\mathbf{{\Theta}}}}^{*}$ and $\Phi:\mathbb{R}\rightarrow[0,1]$ is a known cumulative distribution function (CDF). Given the function $\Phi$ and the matrix ${\bm{\mathbf{{Y}}}}$, the goal of 1-bit matrix completion is to estimate the underlying matrix ${\bm{\mathbf{{\Theta}}}}^{*}$, see for example Davenport et al. (2014).
This problem is a variant of the classic matrix completion problem, in which one directly observes the entries $\theta_{ij}^{*}$ for $(i,j)\in\Omega$. In some applications, one can only observe a stochastic value $y_{ij}$ that is non-linearly related to the underlying value $\theta_{ij}^{*}$. For example, in the Netflix challenge (Koren et al., 2009), the goal was to predict the missing entries of a rating matrix whose observed entries were integers ranging from one to five. In the simplest model for recommendation systems, user responses take on only one of two values: “like” or “dislike”, leading to a 1-bit matrix completion problem. Additional applications involving completion of partially observed binary matrices include quantum state tomography (Gross et al., 2010) and link prediction in network analysis (Liben-Nowell and Kleinberg, 2003).
@@ 1.1 Prior Work
Additionally, the set $\Omega$ of observed entries needs to be well spread-out. Several previous works assumed that $\Omega$ is distributed uniformly at random. Then if $|\Omega|$ is sufficiently large and the above assumptions on ${\bm{\mathbf{{\Theta}}}}^{*}$ hold, it becomes possible to accurately estimate ${\bm{\mathbf{{\Theta}}}}^{*}$, see for example Bhaskar and Javanmard (2015), Davenport et al. (2014), and Cai and Zhou (2013).
Instead of the original problem (2), some works minimized (3) subject to a constraint on the rank and an additional constraint on $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\infty}$ (Bhaskar and Javanmard, 2015; Ni and Gu, 2016). Other works employed a constraint on a convex relaxation of the rank in conjunction with the constraint on $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\infty}$. For example, Davenport et al. (2014) replaced the nonconvex rank constraint with a constraint on the trace norm $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{*}$. They minimized the negative log-likelihood (3) under two constraints $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{*}\leq\alpha\beta\sqrt{mn}$ and $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\infty}\leq\alpha$, where $\alpha$ and $\beta$ are two tuning parameters. Instead of the trace norm, Cai and Zhou (2013) employed the max-norm (Linial et al., 2007) as a convex relaxation of the rank. They minimized the objective (3) under the two constraints $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\max}\leq\alpha\beta$ and $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\infty}\leq\alpha$, where $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\text{max}}=\min_{{\bm{\mathbf{{\Theta}}}}={\bm{\mathbf{{U}}}}{\bm{\mathbf{{V}}}}^{\sf T}}\{\lVert{\bm{\mathbf{{U}}}}\lVert_{2,\infty},\lVert{\bm{\mathbf{{V}}}}\lVert_{2,\infty}\}$ and $\lVert{\bm{\mathbf{{U}}}}\lVert_{2,\infty}$ is the largest $l_{2}$-norm of the rows in ${\bm{\mathbf{{U}}}}$.
Most algorithms for these prior works, however, are non-trivial to implement or computationally intensive. Bhaskar and Javanmard (2015) employ a log-barrier penalty to approximate the infinity norm constraint in combination with a cross-validation procedure to determine how to scale the penalty. The algorithm in Davenport et al. (2014), at each iteration, runs an inner ADMM algorithm to compute the projection onto the intersection of an infinity norm ball and nuclear norm ball. The algorithm in Cai and Zhou (2013), at each iteration, requires projection onto a max-norm constraint which involves solving a nonconvex optimization problem iteratively in conjunction with a rank selection step. Hence some of these methods do not scale well to large matrices and are prohibitively slow. As illustrated later on in Section 4, publicly available implementations may take hours to complete medium-sized matrices.
@@ 1.2 Our Contributions
The main contribution of this work is the derivation of a simple and computationally fast approach to minimize (2). Our iterative method, denoted Majorization-Minimization Gauss-Newton (MMGN), relies on a majorization-minimization strategy. This leads to solving a sequence of standard simpler matrix completion problems. We approximately solve these problems by a single Gauss-Newton step, each of which requires solving a least squares problem. Hence, our method is simple to implement, requiring only standard linear algebra tools. MMGN is typically at least an order of magnitude faster than existing 1-bit matrix completion methods while achieving comparable or even superior estimation accuracy. In addition, MMGN is faster than generic optimization methods that minimize the original objective (2), including gradient descent, quasi-Newton methods, and manifold optimization. Beyond its computational advantages, MMGN’s performance degrades more slowly as the spikiness of the underlying matrix increases compared to alternative methods. In our simulations, as we increased the spikiness of the underlying matrix, the estimation error of other methods increased significantly, whereas MMGN still was able to provide relatively accurate reconstructions.
@@ 1.3 Notation and Organization
We denote vectors by boldface lowercase letters and matrices by boldface capital letters, e.g., ${\bm{\mathbf{{a}}}}\in\mathbb{R}^{n}$ and ${\bm{\mathbf{{A}}}}\in\mathbb{R}^{m\times n}$. We denote the entries of a vector ${\bm{\mathbf{{a}}}}$ and matrix ${\bm{\mathbf{{A}}}}$ by ${a}_{i}$ and ${a}_{ij}$, respectively. The rank of a matrix ${\bm{\mathbf{{A}}}}$ is denoted by $\text{rank}({\bm{\mathbf{{A}}}})$, its Moore–Penrose pseudo inverse by ${\bm{\mathbf{{A}}}}^{\dagger}$, and its column-major vectorization, i.e., the vector obtained by stacking the columns of ${\bm{\mathbf{{A}}}}$ one after the other, by $\text{vec}({\bm{\mathbf{{A}}}})$. We will use the following semi-norm of a matrix ${\bm{\mathbf{{A}}}}$, denoted by $\lVert{\bm{\mathbf{{A}}}}\rVert_{\text{F}(\Omega)}=\sqrt{\sum_{(i,j)\in\Omega}a_{ij}^{2}}$, where $\Omega$ is a subset of ${\bm{\mathbf{{A}}}}$’s indices. Similarly, for a vector ${\bm{\mathbf{{a}}}}$, we denote $\lVert{\bm{\mathbf{{a}}}}\rVert_{\Omega}=\sqrt{\sum_{i\in\Omega}a_{i}^{2}}$, where $\Omega$ is a subset of ${\bm{\mathbf{{a}}}}$’s indices. For two matrices ${\bm{\mathbf{{A}}}}$ and ${\bm{\mathbf{{B}}}}$ of the same size, we denote the Hadamard product as ${\bm{\mathbf{{A}}}}\circ{\bm{\mathbf{{B}}}}$, namely $({\bm{\mathbf{{A}}}}\circ{\bm{\mathbf{{B}}}})_{ij}={a}_{ij}{b}_{ij}$. Similarly, the element-wise quotient of ${\bm{\mathbf{{A}}}}$ and ${\bm{\mathbf{{B}}}}$ is ${\bm{\mathbf{{A}}}}\oslash{\bm{\mathbf{{B}}}}$, namely $({\bm{\mathbf{{A}}}}\oslash{\bm{\mathbf{{B}}}})_{ij}={a}_{ij}/{b}_{ij}$. For a set $S$, we denote its cardinality as $|S|$. Finally, for any univariate function $\Phi:\mathbb{R}\rightarrow\mathbb{R}$, we denote the matrix obtained by applying $\Phi$ entry-wise to ${\bm{\mathbf{{A}}}}$ by $\Phi({\bm{\mathbf{{A}}}})$.
The rest of the paper is organized as follows. Section 2 introduces the MMGN method. In Section 3 we discuss several related problems and approaches similar to ours to solve them. Section 4 presents an empirical evaluation of MMGN with simulations and a real data example. Section 5 concludes with a discussion.
@@ 2 The MMGN Method
We first present an overview of our approach. Given a target rank $r$, our strategy for solving the nonconvex problem (2) is based on the majorization-minimization (MM) principle. Concretely, we replace the objective function $\ell({\bm{\mathbf{{\Theta}}}})$ with a simpler surrogate function, called a majorization. As we show below, minimizing our majorization leads to a standard low-rank matrix completion problem. Hence, MMGN solves a sequence of optimization problems of the form
@@ 2.1 The Majorization-Minimization Principle
The logistic model with $\Phi=1/(1+e^{-x/\sigma})$ satisfies assumptions A1 and A2 with Lipschitz constant $L=\frac{1}{4\sigma^{2}}$. Hence, a majorization is given by Eq. (7) with
@@ 2.2 Inexact Majorization-Minimization via Gauss-Newton
Approximating the nonlinear least squares problem with the linear least squares problem in (13) corresponds to taking a single Gauss-Newton iteration. This approach produces following inexact solution of (10), denoted $\hat{{\bm{\mathbf{{\Theta}}}}}_{\text{MM}}$,
@@
We address three important details about our approach. The first is that the linear least squares problem in (13) has infinitely many solutions. For example, suppose that $(\Delta{\bm{\mathbf{{U}}}}^{*},\Delta{\bm{\mathbf{{V}}}}^{*})$ is a solution to (13), then $(\Delta{\bm{\mathbf{{U}}}}^{*}+{\bm{\tilde{\mathbf{{U}}}}}{\bm{\mathbf{{R}}}},\Delta{\bm{\mathbf{{V}}}}^{*}-{\bm{\tilde{\mathbf{{V}}}}}{\bm{\mathbf{{R}}}}^{\sf T})$ is also a solution to (13) for any ${\bm{\mathbf{{R}}}}\in\mathbb{R}^{r\times r}$. Here we adopt the strategy used by Bauch et al. (2021) and Zilber and Nadler (2022) and select the least $l_{2}$-norm solution to the problem, namely $(\Delta{\bm{\mathbf{{U}}}}^{*},\Delta{\bm{\mathbf{{V}}}}^{*})$ with smallest value of $\lVert\Delta{\bm{\mathbf{{U}}}}^{*}\rVert_{\text{F}}^{2}+\lVert\Delta{\bm{\mathbf{{V}}}}^{*}\rVert_{\text{F}}^{2}$ among all $(\Delta{\bm{\mathbf{{U}}}}^{*},\Delta{\bm{\mathbf{{V}}}}^{*})$ that solve (13). This solution can be computed efficiently using the LSQR algorithm (Paige and Saunders, 1982) or the conjugate gradients method (Hestenes et al., 1952; Kammerer and Nashed, 1972) applied to the normal equations.
@@
The second detail is that the update in (14) may not necessarily lead to a decrease in the objective function. The Gauss-Newton method is an instance of the steepest descent algorithm, and consequently the solution $(\Delta{\bm{\mathbf{{U}}}}^{*},\Delta{\bm{\mathbf{{V}}}}^{*})$ corresponds to a descent direction. The update in (14), however, corresponds to a full Gauss-Newton step which may not necessarily decrease the value of the original objective function in (10). To overcome this potential problem, if the $\hat{{\bm{\mathbf{{\Theta}}}}}_{\text{MM}}$ in (14) does not decrease the original objective, we apply the Armijo backtracking line search to select a suitable stepsize instead of applying a full Gauss-Newton step (Nocedal and Wright, 2006, Chapter 3). The Armijo backtracking line search requires computing the updated factor matrices and the objective value, which costs $\mathcal{O}(|\Omega|r)$ flops. In our experience, however, the backtracking line search is seldom triggered in practice. Details on this and a discussion on the convergence of MMGN are in the supplementary materials.
@@
The third detail is that there is no need to globally minimize the majorization (10). Motivated by computational considerations, we inexactly solve the MM optimization problem by taking a single Gauss-Newton step. This differs from the approach of Zilber and Nadler (2022), which solves (13) using multiple iteration steps and is thus more computationally intensive. Inexact minimization is a standard approach in cases where exact minimization requires an iterative solver. See, for example, the split-feasibility algorithm in Xu et al. (2018), which also employed a single Gauss-Newton step. Under suitable regularity conditions, little is lost by a single-step MM-gradient approach. Not only does it avoid potentially expensive inner iterations within outer MM iterations, but it often exhibits the same local convergence rate as exact minimization (Hunter and Lange, 2004). Algorithm 1 presents the complete MMGN algorithm.

### v3

@@ Abstract
In 1-bit matrix completion, the aim is to estimate an underlying low-rank matrix from a partial set of binary observations. We propose a novel method for 1-bit matrix completion called Majorization-Minimization Gauss-Newton (MMGN). Our method is based on the majorization-minimization principle, which converts the original optimization problem into a sequence of standard low-rank matrix completion problems. We solve each of these sub-problems by a factorization approach that explicitly enforces the assumed low-rank structure and then apply a Gauss-Newton method. Using simulations and a real data example, we illustrate that in comparison to existing 1-bit matrix completion methods, MMGN outputs comparable if not more accurate estimates. In addition, it is often significantly faster, and less sensitive to the spikiness of the underlying matrix. In comparison with three standard generic optimization approaches that directly minimize the original objective, MMGN also exhibits a clear computational advantage, especially when the fraction of observed entries is small.
@@ 1 Introduction
In this work, we consider the following 1-bit matrix completion problem: Let ${\bm{\mathbf{{\Theta}}}}^{*}\in\mathbb{R}^{m\times n}$ be an unknown matrix of exact rank $r^{*}$. The observed data are a subset $\Omega\subset[m]\times[n]$ and a binary matrix ${\bm{\mathbf{{Y}}}}$. At entries $(i,j)\notin\Omega$, $y_{ij}=0$. At entries $(i,j)\in\Omega$, the observations are independent draws as follows,
@@
where ${\theta}_{ij}^{*}$ is the $ij$-th entry of ${\bm{\mathbf{{\Theta}}}}^{*}$, and $\Phi:\mathbb{R}\rightarrow[0,1]$ is a known cumulative distribution function (CDF). Given the function $\Phi$ and the matrix ${\bm{\mathbf{{Y}}}}$, the goal of 1-bit matrix completion is to estimate the underlying matrix ${\bm{\mathbf{{\Theta}}}}^{*}$, see for example Davenport et al. (2014).
This problem is a variant of the standard matrix completion problem, in which one directly observes the entries $\theta_{ij}^{*}$ for $(i,j)\in\Omega$. In some applications, one can only observe a stochastic value $y_{ij}$ that is non-linearly related to the underlying value $\theta_{ij}^{*}$. For example, in the Netflix challenge (Koren et al., 2009), the goal was to predict the missing entries of a rating matrix whose observed entries were integers ranging from one to five. In the simplest model for recommendation systems, user responses take on only one of two values: “like” or “dislike”, leading to a 1-bit matrix completion problem. Additional applications involving completion of partially observed binary matrices include quantum state tomography (Gross et al., 2010) and link prediction in network analysis (Liben-Nowell and Kleinberg, 2003).
@@ 1.1 Prior Work
Additionally, the set $\Omega$ of observed entries needs to be well spread-out. Several previous works assumed that $\Omega$ is distributed uniformly at random. Then if $|\Omega|$ is sufficiently large and the above assumptions on ${\bm{\mathbf{{\Theta}}}}^{*}$ hold, it is possible to accurately estimate ${\bm{\mathbf{{\Theta}}}}^{*}$, see for example Bhaskar and Javanmard (2015), Davenport et al. (2014), and Cai and Zhou (2013).
Instead of the original problem (2), some works minimized (3) subject to a constraint on the rank and an additional constraint on $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\infty}$ (Bhaskar and Javanmard, 2015; Ni and Gu, 2016). Other works employed a constraint on a convex relaxation of the rank in conjunction with the constraint on $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\infty}$. For example, Davenport et al. (2014) replaced the rank by the trace norm $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{*}$. They minimized the negative log-likelihood (3) under two constraints $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{*}\leq\alpha\beta\sqrt{mn}$ and $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\infty}\leq\alpha$, where $\alpha$ and $\beta$ are two tuning parameters. Instead of the trace norm, Cai and Zhou (2013) employed the max-norm (Linial et al., 2007) as a convex relaxation of the rank. They minimized the objective (3) under the two constraints $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\max}\leq\alpha\beta$ and $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\infty}\leq\alpha$, where $\lVert{\bm{\mathbf{{\Theta}}}}\lVert_{\text{max}}=\min_{{\bm{\mathbf{{\Theta}}}}={\bm{\mathbf{{U}}}}{\bm{\mathbf{{V}}}}^{\sf T}}\{\lVert{\bm{\mathbf{{U}}}}\lVert_{2,\infty},\lVert{\bm{\mathbf{{V}}}}\lVert_{2,\infty}\}$ and $\lVert{\bm{\mathbf{{U}}}}\lVert_{2,\infty}$ is the largest $l_{2}$-norm of the rows in ${\bm{\mathbf{{U}}}}$.
Most algorithms for these prior works, however, are computationally intensive and non-trivial to implement. Bhaskar and Javanmard (2015) employed a log-barrier penalty to approximate the infinity norm constraint in combination with a cross-validation procedure to determine how to scale the penalty. The algorithm in Davenport et al. (2014), at each iteration, runs an inner ADMM algorithm to compute the projection onto the intersection of an infinity norm ball and nuclear norm ball. The algorithm in Cai and Zhou (2013), at each iteration, requires projection onto a max-norm constraint which involves solving a nonconvex optimization problem iteratively in conjunction with a rank selection step. Hence some of these methods do not scale well to large matrices and are prohibitively slow. As illustrated later on in Section 4, publicly available implementations may take hours to complete medium-sized matrices.
@@ 1.2 Our Contributions
The main contribution of this work is the derivation of a simple and computationally fast approach to minimize (2). Our iterative method, denoted Majorization-Minimization Gauss-Newton (MMGN), relies on a majorization-minimization strategy. This leads to solving a sequence of standard simpler matrix completion problems. We approximately solve these problems by a single Gauss-Newton step, each of which requires solving a least squares problem. Hence, our method is simple to implement, requiring only standard linear algebra tools. MMGN is typically at least an order of magnitude faster than existing 1-bit matrix completion methods while achieving comparable or even superior estimation accuracy. In addition, MMGN is faster than generic optimization methods that minimize the original objective (2), including gradient descent, quasi-Newton methods, and manifold optimization algorithms. Beyond its computational advantages, as illustrated via simulations, MMGN’s performance degrades more slowly as the spikiness of the underlying matrix increases compared to alternative methods.
@@ 1.3 Notation and Organization
We denote vectors by boldface lowercase letters and matrices by boldface capital letters, e.g., ${\bm{\mathbf{{a}}}}\in\mathbb{R}^{n}$ and ${\bm{\mathbf{{A}}}}\in\mathbb{R}^{m\times n}$. We denote the entries of a vector ${\bm{\mathbf{{a}}}}$ and matrix ${\bm{\mathbf{{A}}}}$ by ${a}_{i}$ and ${a}_{ij}$, respectively. The rank of a matrix ${\bm{\mathbf{{A}}}}$ is denoted by $\text{rank}({\bm{\mathbf{{A}}}})$, its Moore–Penrose pseudo inverse by ${\bm{\mathbf{{A}}}}^{\dagger}$, and its column-major vectorization, i.e., the vector obtained by stacking the columns of ${\bm{\mathbf{{A}}}}$ one after the other, by $\text{vec}({\bm{\mathbf{{A}}}})$. We will use the following semi-norm of a matrix ${\bm{\mathbf{{A}}}}$, denoted by $\lVert{\bm{\mathbf{{A}}}}\rVert_{\text{F}(\Omega)}=\sqrt{\sum_{(i,j)\in\Omega}a_{ij}^{2}}$, where $\Omega$ is a subset of ${\bm{\mathbf{{A}}}}$’s indices. Similarly, for a vector ${\bm{\mathbf{{a}}}}$, we denote $\lVert{\bm{\mathbf{{a}}}}\rVert_{\Omega}=\sqrt{\sum_{i\in\Omega}a_{i}^{2}}$, where $\Omega$ is a subset of ${\bm{\mathbf{{a}}}}$’s indices. For two matrices ${\bm{\mathbf{{A}}}}$ and ${\bm{\mathbf{{B}}}}$ of the same size, we denote the Hadamard product by ${\bm{\mathbf{{A}}}}\circ{\bm{\mathbf{{B}}}}$, namely $({\bm{\mathbf{{A}}}}\circ{\bm{\mathbf{{B}}}})_{ij}={a}_{ij}{b}_{ij}$. Similarly, the element-wise quotient of ${\bm{\mathbf{{A}}}}$ and ${\bm{\mathbf{{B}}}}$ is ${\bm{\mathbf{{A}}}}\oslash{\bm{\mathbf{{B}}}}$, namely $({\bm{\mathbf{{A}}}}\oslash{\bm{\mathbf{{B}}}})_{ij}={a}_{ij}/{b}_{ij}$. For a set $S$, we denote its cardinality by $|S|$. Finally, for any univariate function $\Phi:\mathbb{R}\rightarrow\mathbb{R}$, we denote the matrix obtained by applying $\Phi$ entry-wise to ${\bm{\mathbf{{A}}}}$ by $\Phi({\bm{\mathbf{{A}}}})$.
The rest of the paper is organized as follows. Section 2 introduces the MMGN method. Section 3 discusses several related problems and approaches similar to ours to solve them. Section 4 presents an empirical evaluation of MMGN with simulations and a real data example. Section 5 concludes with a discussion.
@@ 2 The MMGN Method
We first present an overview of our approach. Given a target rank $r$, our strategy for solving the nonconvex problem (2) is based on the majorization-minimization (MM) principle. Concretely, we replace the objective $\ell({\bm{\mathbf{{\Theta}}}})$ with a simpler surrogate function, called a majorization. As we show below, minimizing our majorization leads to a standard low-rank matrix completion problem. Hence, MMGN solves a sequence of optimization problems of the form
@@ 2.1 The Majorization-Minimization Principle
The logistic model with $\Phi(x)=1/(1+e^{-x/\sigma})$ satisfies assumptions A1 and A2 with Lipschitz constant $L=\frac{1}{4\sigma^{2}}$. Hence, a majorization is given by Eq. (7) with
@@ 2.2 Inexact Majorization-Minimization via Gauss-Newton
Approximating the nonlinear least squares problem with the linear least squares problem in (13) corresponds to taking a single Gauss-Newton iteration. This approach produces the following inexact solution of (10), denoted $\hat{{\bm{\mathbf{{\Theta}}}}}_{\text{MM}}$,
@@
We address three important details about our approach. First, the linear least squares problem (13) has infinitely many solutions. For example, suppose that $(\Delta{\bm{\mathbf{{U}}}}^{*},\Delta{\bm{\mathbf{{V}}}}^{*})$ is a solution to (13), then $(\Delta{\bm{\mathbf{{U}}}}^{*}+{\bm{\tilde{\mathbf{{U}}}}}{\bm{\mathbf{{R}}}},\Delta{\bm{\mathbf{{V}}}}^{*}-{\bm{\tilde{\mathbf{{V}}}}}{\bm{\mathbf{{R}}}}^{\sf T})$ is also a solution for any ${\bm{\mathbf{{R}}}}\in\mathbb{R}^{r\times r}$. Here we adopt the strategy used by Bauch et al. (2021) and Zilber and Nadler (2022) and select the least $l_{2}$-norm solution, namely $(\Delta{\bm{\mathbf{{U}}}}^{*},\Delta{\bm{\mathbf{{V}}}}^{*})$ with smallest value of $\lVert\Delta{\bm{\mathbf{{U}}}}^{*}\rVert_{\text{F}}^{2}+\lVert\Delta{\bm{\mathbf{{V}}}}^{*}\rVert_{\text{F}}^{2}$ among all $(\Delta{\bm{\mathbf{{U}}}}^{*},\Delta{\bm{\mathbf{{V}}}}^{*})$ that solve (13). This solution can be computed efficiently using the LSQR algorithm (Paige and Saunders, 1982) or the conjugate gradient method (Hestenes et al., 1952; Kammerer and Nashed, 1972) applied to the normal equations.
@@
The second detail is that the update in (14) may not necessarily lead to a decrease in the objective function. The Gauss-Newton method is an instance of the steepest descent algorithm, and consequently the solution $(\Delta{\bm{\mathbf{{U}}}}^{*},\Delta{\bm{\mathbf{{V}}}}^{*})$ corresponds to a descent direction. In fact, we prove in the supplementary materials that any solution to (13) is a descent direction. The update in (14), however, corresponds to a full Gauss-Newton step which may not necessarily decrease the value of the original objective function in (10). To overcome this potential problem, if $\hat{{\bm{\mathbf{{\Theta}}}}}_{\text{MM}}$ in (14) does not decrease the original objective, we apply the Armijo backtracking line search to select a suitable stepsize (Nocedal and Wright, 2006, Chapter 3). The Armijo backtracking line search requires computing the objective value, which costs $\mathcal{O}(|\Omega|r)$ flops. In our experience, however, the backtracking line search is seldom triggered in practice. Details on this and a discussion on the convergence of MMGN are in the supplementary materials.
@@
The third detail is that there is no need to globally minimize the majorization (10). Motivated by computational considerations, we inexactly solve the MM optimization problem by taking a single Gauss-Newton step. This differs from the approach of Zilber and Nadler (2022), which solves (13) using multiple iteration steps and is thus more computationally intensive. Inexact minimization is a standard approach in cases where exact minimization requires an iterative solver. See, for example, the split-feasibility algorithm in Xu et al. (2018), which also employed a single Gauss-Newton step. Under suitable regularity conditions, little is lost by a single-step MM-gradient approach. Not only does it avoid potentially expensive inner iterations within outer MM iterations, but it often exhibits the same local convergence rate as exact minimization (Hunter and Lange, 2004).
@@
We present the complete MMGN algorithm in Algorithm 1. MMGN requires $\mathcal{O}(|\Omega|r)$ operations per iteration, which is comparable to the gradient-based methods. Details on these per-iteration costs can be found in Section C of the supplement. Empirically, however, MMGN consistently outperforms generic gradient-based methods in terms of runtime as shown in Section 4. We close this section with a brief discussion to understand why this might be expected.
For simplicity, we work with a bivariate instance of the objective function for the factorization approach. Let $\ell:\mathbb{R}\mapsto\mathbb{R}$ be twice differentiable with an $L$-Lipschitz first derivative. Define $f:\mathbb{R}\times\mathbb{R}\mapsto\mathbb{R}$ as $f(u,v)=\ell(uv)$. Let $u=\tilde{u}+\Delta u$ and $v=\tilde{v}+\Delta v$ and consider the second order Taylor expansion of $f$ around $(\tilde{u},\tilde{v})$.
$f(u,v)$ $=$ $\ell(\tilde{u}\tilde{v})+\ell^{\prime}(\tilde{u}\tilde{v})[\Delta u\tilde{v}+\tilde{u}\Delta v]+\frac{\ell^{\prime\prime}(\tilde{u}\tilde{v})}{2}[\Delta u\tilde{v}+\tilde{u}\Delta v]^{2}+r(\Delta u,\Delta v),$ (15)
where $r(\Delta u,\Delta v)$ is $o((\Delta u+\Delta v)^{3})$.
MMGN employs the following majorization of $f(u,v)$ at $(\tilde{u},\tilde{v})$
$g(u,v\mid\tilde{u},\tilde{v})$ $=$ $\ell(\tilde{u}\tilde{v})+\ell^{\prime}(\tilde{u}\tilde{v})[\Delta u\tilde{v}+\tilde{u}\Delta v]+\frac{L}{2}[\Delta u\tilde{v}+\tilde{u}\Delta v]^{2}.$ (16)
There are two differences between (15) and (16). First, MMGN uses the uniform bound $L$ in lieu of computing the second derivative ′′$\ell^{\prime\prime}(\tilde{u}\tilde{v})$. This substitution greatly simplifies computations in the general case when $\ell$ is a function of matrices ${\bm{\mathbf{{U}}}}\in\mathbb{R}^{m\times r}$ and ${\bm{\mathbf{{V}}}}\in\mathbb{R}^{n\times r}$. Second, MMGN drops the third order error term $r(\Delta u,\Delta v)$. Recall that the minimizer to (16) is not unique and that MMGN chooses the least 2-norm minimizer. This has the beneficial side-effect of finding a descent direction along which the error term $r(\Delta u,\Delta v)$ is small. In other words, MMGN picks a descent direction, among many possible descent directions, along which the surrogate function has a small approximation error. This small error is likely the reason the full Gauss-Newton step is almost always taken by MMGN.
By contrast, the gradient descent surrogate is
$h(u,v\mid\tilde{u},\tilde{v})$ $=$ $\ell(\tilde{u}\tilde{v})+\ell^{\prime}(\tilde{u}\tilde{v})[\Delta u\tilde{v}+\tilde{u}\Delta v]+\frac{1}{2\alpha}[\Delta u^{2}+\Delta v^{2}],$ (17)
where $\alpha$ is the step-size parameter.
There are again two differences to note. First, (17) uses the isotropic quadratic function
$\frac{1}{2\alpha}(\Delta u^{2}+\Delta v^{2})$
to approximate the anisotropic quadratic function
′′$\frac{\ell^{\prime\prime}(\tilde{u}\tilde{v})}{2}(\Delta u\tilde{v}+\tilde{u}\Delta v)^{2}$
in (15). Moreover, the quadratic term in (17) does not even depend on the current iterate $(\tilde{u},\tilde{v})$. By contrast, MMGN uses a similar anisotropic quadratic function in (16) as in (15). The quadratic dependency in (16) on $(\tilde{u},\tilde{v})$ and $(\Delta u,\Delta v)$ is the same up to a scalar multiple as in (15).
Second, the update direction $(\Delta u,\Delta v)$ chosen by gradient descent can be large when far from a solution. Thus, $r(\Delta u,\Delta v)$ could be large and so the surrogate approximation in (17) may be poor along the gradient direction. This is a well known issue with gradient descent that typically results in gradient descent taking very small steps when the objective function has steep and narrow valleys.
Now recall that the per-iteration cost of MMGN is $O(\lvert\Omega\rvert r)$. The cost of computing the gradient of $\ell({\bm{\mathbf{{U}}}}{\bm{\mathbf{{V}}}}^{\sf T})$ is also $O(\lvert\Omega\rvert r)$. The above simple analysis suggests that MMGN computes updates from surrogate models that capture useful second order information with a per-iteration cost that is roughly the same cost as a gradient evaluation. In other words, we expect MMGN to behave as a quasi-Newton method that has a per-iteration cost similar to a first order method.